OpenStudy (anonymous):

An object is dropped from a height of 25 meters. At what velocity will it hit the ground? A. 7.0 meters/second B. 11 meters/second C. 22 meters/second D. 49 meters/second E. 70 meters/second

OpenStudy (souvik):

use this equation..v^2=u^2+2gh

OpenStudy (anonymous):

An engine moves a motorboat through water at a constant velocity of 22 meters/second. If the force exerted by the motor on the boat is 5.75 × 103 newtons, what is the power of the motor? A. 7.9 × 104 watts B. 1.3 × 105 watts C. 2.5 × 105 watts D. 3.7 × 105 watts E. 5.0 × 105 watts

OpenStudy (souvik):

power P=F*v

OpenStudy (anonymous):

A skateboarder with a mass of 45 kilograms is riding on a skateboard with a mass of 2.5 kilograms. What should be the velocity of the skateboarder if she has to attain a momentum of 264 kilogram meters/second? A. 5.5 meters/second B. 55.5 meters/second C. 1.2 × 102 meters/second D. 1.2 × 103 meters/second E. 1.2 × 104 meters/second

OpenStudy (anonymous):

do you know the answer?

OpenStudy (souvik):

total mass is 45+2.5=47.5 kg m*v=264 find v

OpenStudy (anonymous):

A 55 newton force applied on an object moves the object 10 meters in the same direction as the force. What is the value of work done on the object? A. 0 joules B. 5.5 joules C. -5.5 joules D. 5.5 × 102 joules E. -5.5 × 102 joules

OpenStudy (souvik):

W=F*d

OpenStudy (anonymous):

b or d? :O cause I get 550 but theres nothing close to that

OpenStudy (souvik):

550=5.5*10^2

OpenStudy (anonymous):

nvm found it(:

OpenStudy (anonymous):

An object of mass weighing 5.24 kilograms is raised to a height of 1.63 meters. What is the potential energy of the object at that height? Note that acceleration due to gravity is 9.8 meters/second2? A. 65 joules B. 84 joules C. 91 joules D. 1.0 × 102 joules E. 1.5 × 102 joules

OpenStudy (souvik):

E=mgh

OpenStudy (anonymous):

you make it seem so easyyyy (:

OpenStudy (anonymous):

The landing of a spacecraft is cushioned with the help of airbags. During its landing on Mars, the velocity of downward fall is 16 meters/second. Immediately after the impact, the velocity is reduced to 1.2 meters/second. If the spacecraft has a mass of 11.5 × 104 kilograms, what is your estimate for the impulse if the time of impact is 0.8 seconds? A. 7.5 × 105 kilogram meters/second B. 9.0 × 106 kilogram meters/second C. 9.4 × 104 kilogram meters/second D. 2.1 × 106 kilogram meters/second E. 1.0 × 103 kilogram meters/second

OpenStudy (souvik):

tell me what is impulse..?

OpenStudy (anonymous):

it is the integral of a force with respect to time, which gives you the change in the momentum of the body being acted on by the force.

OpenStudy (souvik):

so calculate the change in momentum first...

OpenStudy (anonymous):

idk how?/:

OpenStudy (souvik):

okay first find change in velocity..

OpenStudy (anonymous):

15.8?

OpenStudy (souvik):

well done...now find change in momentum..

OpenStudy (anonymous):

how do you find that

OpenStudy (souvik):

mass*change in velocity..

OpenStudy (theeric):

Haha, @souvik, you deserve like 10 medals!

OpenStudy (souvik):

haha..@theEric

OpenStudy (anonymous):

yes he does more than 10 medals!!!!!!! :)